# Towards End-to-End Neuromorphic Voxel-based 3D Object Reconstruction Without Physical Priors

### v1

@@ Abstract
Neuromorphic Camera, 3D Reconstruction, Deep Learning
@@ I Introduction
3D reconstruction in VR/AR applications enables realistic restoration of scenes and objects, serving as a 3D form of information present that provides users with a more immersive experience [1, 2]. Many devices can be used to collect data for 3D reconstruction, including traditional RGB cameras, RGB-D cameras, LiDAR, structured light systems, etc. However, they have different limitations, including limited dynamic range, motion blur, high power consumption, etc. [3].
@@
Neuromorphic cameras in 3D reconstruction tasks can be divided into stereo and monocular types. Stereo neuromorphic cameras involve multiple rigidly connected neuromorphic cameras, providing information from multiple viewpoints. These methods typically perform scene scanning and produce real-time semi-dense reconstruction results. They often follow the classical two-step stereo solution: matching disparities at the same timestamps and then triangulating 3D points to calculate the depth information of every point in the scene [6, 3]. However, performing 3D reconstruction with a monocular neuromorphic camera, which lacks disparity information, often requires more complex computations. Such methods must be combined with the physical prior knowledge (e.g., camera trajectories) to achieve similar results as parallax in the stereo task. The physical prior is always required, and it can either be obtained by simultaneous mapping with other devices or be predicted by a Visual Odometer (VO) or SLAM algorithm targeting the event stream [6]. Moreover, these methods all require a complex event-to-3D pipeline [7, 8, 9, 10]. The pipeline is a full workflow that optimally processes event data and reconstructs it into a 3D model, including steps such as event representation, prior estimation, computation of parallax, triangulation, depth estimation, and 3D model reconstruction. Event processing pipelines are a common research direction in these tasks. A recent study introduced a neural network, E2V [1], capable of directly taking the represented event stream as input and outputting voxel results. This approach made progress in simplifying the event processing pipeline. It is also the first method to perform voxel-based 3D reconstruction using a monocular neuromorphic camera. However, it leaves substantial room for improvement in reconstruction accuracy.
@@ II-B Deep Learning-based Methods
Recent methods using monocular neuromorphic cameras for 3D reconstruction rely on existing synthetic or real event datasets to train deep learning models for dense 3D reconstruction results. However, pipelines and the estimation of physical priors remain indispensable.
@@ III-A Novel Event Representation: Sobel Event Frame
Event representation refers to preprocessing the data captured by the neuromorphic camera, which encodes the information of the event’s coordinate, timestamp, and polarity.
Event Frame is the most commonly used event representation method, referring to all approaches that can frame event streams. Event Frame divides event data into multiple fixed-length time windows, either based on timestamp or event count. Each time window generates an image-like frame where each pixel represents whether an event occurred within that time window. Pixels where no event occurred during the time window are filled with 0.
@@
The Sobel operator is an image processing algorithm primarily used for edge detection. It computes the gradient of pixel intensity in both the horizontal (x-axis) and vertical (y-axis) directions using convolution, aiming to highlight the edges in the image [15].
In neuromorphic camera-related research, there has been some work on applying convolution to event data, but only a few studies have used the Sobel operator [16, 17], and these are applied on initial event stream aimming to detect edge and structure in dynamic scenes, which can be considered as non-frame-based preprocessing method for events.
@@ III-B Event-to-Voxel 3D ResNet with ECA
The E2V model uses a deeper model architecture, ResNet-152, as the encoder to extract complex 3D features from the represented event stream [1]. However, due to massive information in 3D data, the extracted features may include irrelevant or redundant information (e.g., the object is primarily located in the center of the event frame, while no events occur around the edges), which limits the performance of model.
@@
For decoder, we add a controllable output reshaping module to E2V decoder [1], learning the correlation between the event data and ground truth labels, to generate the voxel logit output.
@@ III-C Optimal Binarization Threshold Selection Principle
Output Binary $\text{Output Binary Value}=\begin{cases}1,&\text{if }\sigma(x_{\text{out}})>p,\\ 0,&\text{if }\sigma(x_{\text{out}})\leq p.\end{cases}$

### v2

@@ Abstract
Event Camera, 3D Reconstruction, Neuromorphic Vision, Deep Learning
footnotetext: Corresponding author. Email: chxu4146@uni.sydney.edu.au
@@ I Introduction
3D reconstruction in VR/AR applications enables realistic restoration of scenes and objects, serving as a 3D form of information present that provides users with a more immersive experience. Many devices can be used to collect data for 3D reconstruction, including traditional RGB cameras, RGB-D cameras, LiDAR, structured light systems, etc. However, they have different limitations, including limited dynamic range, motion blur, high power consumption, etc. [1].
@@
Neuromorphic cameras in 3D reconstruction tasks can be divided into stereo and monocular types [4]. Stereo neuromorphic cameras involve multiple rigidly connected neuromorphic cameras, providing information from multiple viewpoints. These methods typically perform scene scanning and produce real-time semi-dense reconstruction results. They often follow the classical two-step stereo solution: matching disparities at the same timestamps and then triangulating 3D points to calculate the depth information of every point in the scene [5, 1]. However, performing 3D reconstruction with a monocular neuromorphic camera, which lacks disparity information, often requires more complex computations. Such methods must be combined with the physical prior knowledge (e.g., camera trajectories) to achieve similar results as parallax in the stereo task. The physical prior is always required, and it can either be obtained by simultaneous mapping with other devices or be predicted by a Visual Odometer (VO) or SLAM algorithm targeting the event stream [5]. Moreover, these methods all require a complex event-to-3D pipeline [6, 7, 8, 9]. The pipeline is a full workflow that optimally processes event data and reconstructs it into a 3D model, including steps such as event representation, prior estimation, computation of parallax, triangulation, depth estimation, and 3D model reconstruction [4]. Event processing pipelines are a common research direction in these tasks. A recent study introduced a neural network, E2V [10], capable of directly taking the represented event stream as input and outputting voxel results. This approach made progress in simplifying the event processing pipeline. It is also the first method to perform voxel-based 3D reconstruction using a monocular neuromorphic camera. However, it leaves substantial room for improvement in reconstruction accuracy.
@@ II-B Deep Learning-based Methods
Recent methods using monocular neuromorphic cameras for 3D reconstruction rely on existing synthetic or real event datasets to train deep learning models for dense 3D reconstruction results. However, pipelines and the estimation of physical priors remain indispensable [4].
@@ III-A Novel Event Representation: Sobel Event Frame
Event representation refers to preprocessing the event data, which encodes the event’s coordinate, timestamp, and polarity, aiming for better feature extraction [15, 16].
Event Frame is the most commonly used event representation method [17], referring to all approaches that can frame event streams. Event Frame divides event data into multiple fixed-length time windows, either based on timestamp or event count. Each time window generates an image-like frame where each pixel represents whether an event occurred within that time window. Pixels where no event occurred during the time window are filled with 0.
@@
The Sobel operator is an image-processing algorithm primarily used for edge detection. It computes the gradient of pixel intensity in both the horizontal (x-axis) and vertical (y-axis) directions using convolution, aiming to highlight the edges in the image [18].
In neuromorphic camera-related research, there has been some work on applying convolution to event data, but only a few studies have used the Sobel operator [19, 20], and these are applied on initial event stream aiming to detect edge and structure in dynamic scenes, which can be considered as a non-frame-based preprocessing method for events.
@@ III-B Event-to-Voxel 3D ResNet with ECA
The E2V model uses a deeper model architecture, ResNet-152, as the encoder to extract complex 3D features from the represented event stream [10]. However, due to massive information in 3D data, the extracted features may include irrelevant or redundant information (e.g., the object is primarily located in the center of the event frame, while no events occur around the edges), which limits the performance of the model.
@@
For the decoder, we add a controllable output reshaping module to E2V decoder [10], learning the correlation between the event data and ground truth labels, to generate the voxel logit output.
@@ III-C Optimal Binarization Threshold Selection Principle
Output Binary $\text{Output Binary Value}=\begin{cases}1,&\text{if }\sigma(x_{\text{out}})>p,\\ 0,&\text{if }\sigma(x_{\text{out}})\leq p.\end{cases}$ (2)